Find the volumes in L of the 300 mM MgSO4 solution, 0.2 M NaH2PO4 and glycerol, that you would use to prepare 500 mL of 0.02 M phosphate buffer (pH 7.2), 0.5 M NaCl, 3 mM MgSO4, 1% glycerol (v/v)?

(from pH table for sodium phosphate buffer at 25°C) pH 7.2 --> 28% monobasic, 72% dibasic.
Answer #1

The volumes are calculated:

V MgSO4 = C2 * V2 / C1 = 3 mM * 500 mL / 300 mM = 5 mL

V NaH2PO4 = 0.02 M * 500 mL / 0.2 M * (72/100) = 36 mL

V Glycerol = (1 mL Glycerol / 100 mL solution) * 500 mL solution = 5 mL

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.